OpenStudy (anonymous):

solve for x: log x+log(x-9)=log(20x)

OpenStudy (kirbykirby):

\[\log x + \log(x-9)=\log(20x) \\ \log x+\log(x-9)=\log(20)+\log x\\ \log(x-9)=\log(20)\]

OpenStudy (kirbykirby):

I hope the rest will be obvious :)

OpenStudy (kirbykirby):

The second line, I just used the rule \(\log(ab)=\log(a)+\log(b)\) on log(20x)

OpenStudy (anonymous):

would you bring the log(20) over and make it log[(x-9)/20] ?

OpenStudy (kirbykirby):

Well since you have a log on both sides, you can just "cancel" them out and get x-9 = 20 :)

OpenStudy (anonymous):

oh, i see. that makes sense. thank you! (:

OpenStudy (kirbykirby):

yw :)